Find the magnitude of an electron's electric field if the field itself is 50cm away from it.


1
Expert's answer
2022-03-07T11:43:40-0500

he electric field due the point charge


"E=k\frac{q}{r^2}""E=9*10^9*\frac{1.6*10^{-19}}{0.5^2}=5.76*10^{-9}\:\rm V/m"